Consider the neutralization reaction


2HNO3(aq)+Ba(OH)2(aq)⟶2H2O(l)+Ba(NO3)2(aq)

A  0.110 L sample of an unknown HNO3 solution required 51.1 mL of 0.150 M Ba(OH)2 of

0.150 M Ba(OH)2 for complete neutralization. What is the concentration of the 

HNO3 solution?


1
Expert's answer
2022-04-11T22:09:04-0400

2HNO3(aq) + Ba(OH)2(aq) ⟶ 2H2O(l) + Ba(NO3)2(aq)

n(Ba(OH)2) = C(sol) × V(sol) = 0.150 mol/L × 0.0511 L = 0.007665 mol

2 mol HNO3 - 1 mol Ba(OH)2

x mol HNO3 - 0.007665 mol Ba(OH)2

x = 0.007665 × 2 = 0.01533 mol

C(sol) = n(HNO3) / V(sol) = 0.01533 mol / 0.110 L = 0.139 mol/L